Calculate distances that one vehicle of the organisation travels per day,per week and per month.if you do not work directly with vehicles, calculate the distance you travel to work and back. Show calculations and state the distance in the correct s1 unit.



1
Expert's answer
2022-06-08T01:43:44-0400

Solution:

I do not have a car and travel to work and back by walking. My office is 2000 m away from home. So, "d1=2000m*2=4000m" is the total distance I covered to work and back. Also, everyday in the office I use at least 3 times WC, which downstairs and away from my working room 30 m, and 1 time lunch, which 40 m away, walking inside office about 200 m. So,

"d2=30m*2*3+40m*2+200m=460m;" total distance covered inside the office in one day.

"d(day)=d1+d2=4000m+460m=4460m;"

I work 5 days per week:

"d(week)=5*d(day)=5*4460m=22300m;"

and we have about 22 working days per month:

"d(month)=22*d(week)=22*22300m=490600m."

Answer:

"d(day)=4460m;"

"d(week)=22300m;"

"d(month)=490600m."